OpenStudy (dopeeaf):

Help.? Medal & fan .! A line has a slope of 2 and passes through the point (4,0). What are the coordinates of another point that is on the line? (0,2) (2,0) (6,4) (4,6)

rishavraj (rishavraj):

u see \[\frac{ y_2 - y_1 }{ x_2 - x_1 } = m\] m is the slope

OpenStudy (dopeeaf):

Okay

rishavraj (rishavraj):

@dopeeaf u can proceed further??

jabez177 (jabez177):

http://prntscr.com/9q2okw Say that's the slope of 2 and can you find another point that is on that line?

OpenStudy (dopeeaf):

Well I know/I think that it can't be (6,4) or (4,6)

OpenStudy (anonymous):

with slope, a slope of 2 is rise over run, or, \[\frac{ rise}{ run}\] which makes the slope 2 up and 2 across. graph the equation and you will get your answer, use transum to graph it easier.

OpenStudy (anonymous):

I believe It's A

OpenStudy (anonymous):

excuse me, 2 up and 1 across

OpenStudy (dopeeaf):

Hmmmm, okay thanks !